# Resonant Photon-Axion Mixing Driven by Dark Matter Oscillations

### v1

@@ Abstract
Wave propagation in a coherently oscillating background is intrinsically a periodically driven problem. We show that photon propagation through an axion dark matter background in the presence of a magnetic field is governed by Floquet physics, distinct from conventional static or adiabatic mixing paradigms. Coherent photon–axion mode conversion occurs when the mismatch between the photon and axion dispersion relations is compensated by integer harmonics of the axion oscillation frequency, $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$ ($n\in\mathbb{Z}$) with the axion mass $m_{a}$, even far from the standard level-crossing condition $\Delta_{\gamma}\approx\Delta_{a}$. Crucially, this resonance disappears entirely if the axion oscillations are averaged over, and is therefore systematically missed in conventional static or adiabatic treatments. This driven resonance represents a unitary Floquet mode-mixing process and is fundamentally distinct from parametric instability or stimulated axion decay, preserving the axion dark matter number density. We develop a general Floquet framework for photon propagation in oscillating axion backgrounds, revealing that resonant mixing generates robust polarization effects during propagation. As an astrophysical application, we apply this mechanism to the realistic environment of the blazar 3C 279 to derive concrete constraints on the axion–photon coupling. While the observational manifestation depends on environmental conditions, the underlying driven mixing mechanism is generic to coherent axion dark matter, revealing a previously overlooked regime of photon–axion conversion.
@@
Introduction.– The propagation of waves in time-dependent media is a central problem across physics. When the properties of a medium vary periodically in time, the resulting dynamics are generically governed by Floquet physics, leading to resonant mode mixing that cannot be captured by static or adiabatic approximations. Such driven propagation phenomena are well known in condensed matter and quantum optics [1965PhRv..138..979S, Sambe:1973cnm, 2004PhR...390....1C, Oka:2009djg, Eckardt:2016lof], yet they have received little attention in the context of particle astrophysics.
@@
A prominent and well-motivated realization of a coherently oscillating medium arises if dark matter (DM) [Zwicky:1933gu, Bertone:2018krk] is composed of axions [Peccei:1977hh, Peccei:1977ur, Weinberg:1977ma, Wilczek:1977pj, Cheng:1987gp] or axion-like particles [Preskill:1982cy, Abbott:1982af, Dine:1982ah, Arvanitaki:2009fg, Graham:2015ouw, Peccei:2006as, Marsh:2015xka, DiLuzio:2020wdo, Adams:2022pbo, OHare:2024nmr]. In this case, the DM forms a classical field that oscillates at a frequency set by the axion mass over macroscopic coherence scales. Photons propagating through this background therefore experience a medium whose optical properties are intrinsically periodic in time. Nevertheless, conventional treatments of photon–axion mixing [Raffelt:1987im, Raffelt:1996wa] typically average over the axion oscillations or assume an effectively static background, thereby reducing the problem to a time-independent level-crossing or adiabatic conversion scenario.
In astrophysical environments, the photon–axion system is typically analyzed in two regimes: photon–axion conversion (the Primakoff effect) [Primakoff:1951iae, Sikivie:1983ip, Raffelt:1987im] and rotation of the polarization plane (birefringence) [Carroll:1989vb, Carroll:1998zi, Harari:1992ea]. Conventional analyses employ adiabatic or perturbative approximations, effectively averaging over rapid axion oscillations or treating the background as quasi-static.
@@
In this work, we show that these assumptions miss a qualitatively novel class of resonant propagation effects. We demonstrate that photon–axion mixing in a coherent axion DM background is fundamentally a periodically driven problem, governed by Floquet dynamics rather than static dispersion matching. Our analysis reveals that the intrinsic time periodicity of the axion background acts as a parametric driver, actively inducing coherent mixing between photon and axion modes. This process leads to parametric resonance when the mismatch between the photon and axion dispersion relations is compensated by integer harmonics of the axion mass as $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$ ($n\in\mathbb{Z}$). Consequently, resonant conversion can occur far from the conventional level-crossing resonance condition $\Delta_{\gamma}\approx\Delta_{a}$ [Pshirkov:2007st, Hook:2018iia]. This resonance arises purely from temporal phase matching induced by the oscillating background and does not rely on spatial inhomogeneities or fine-tuned environmental profiles. The driven resonance is not a higher-order correction to static mixing: it relies on the coherent axion oscillation and vanishes completely once this time dependence is averaged out.
The driven resonance identified here is fundamentally distinct from parametric instability or stimulated axion decay ($m_{a}\simeq 2\omega$) [Tkachev:2014dpa, Arza:2019nta, Caputo:2018vmy, Caputo:2018ljp]. The propagation dynamics remain unitary, the axion background number density is conserved, and the system exhibits coherent Floquet mode mixing rather than exponential amplification.
@@
We develop a general Floquet framework for photon propagation in oscillating axion backgrounds and show that this mechanism leads to robust and generic consequences for photon polarization during propagation. While specific observational manifestations depend on environmental conditions, the underlying driven mixing phenomenon is intrinsic to coherent axion DM and represents a previously overlooked regime of photon–axion conversion. Our results establish a systematic framework for analyzing propagation effects in time-dependent backgrounds, revealing previously unexplored resonant regimes.
@@
Driven photon–axion Mixing.– Consider photon propagation in the presence of a magnetic field and a coherently oscillating axion background (see also [Espriu:2014lma, Masaki:2019ggg, McDonald:2019wou] for a similar system),
@@
where $m_{a}$ is the axion mass, $\rho_{\text{DM}}$ is the DM density, and $\varphi_{a}$ is the phase of the axion background. In an appropriate polarization basis, the propagation equation takes the form
@@
where the state vector $\Psi$ includes the two photon polarization states and the axion field. The photon–axion system is governed by a Hermitian mixing Hamiltonian whose coefficients contain an explicit periodic modulation at frequency $m_{a}$, induced by the axion background. $\mathcal{H}_{0}$ incorporates the usual photon dispersion $\Delta_{\perp}$ and $\Delta_{\parallel}$, axion mass term $\Delta_{a}$, and magnetic-field-induced mixing $\Delta_{\mathrm{B}}$, and is given by
@@
$\mathcal{H}_{1}$ encodes the axion-induced birefringence
@@
with
@@
In the linear polarization basis, the axion-induced Faraday term is purely imaginary, reflecting its interpretation as a rotation generator. The Hamiltonian is manifestly Hermitian, and its only explicit time dependence arises from the oscillating axion background:
@@
where the amplitude is $\Delta_{\text{DM}}\equiv g_{a\gamma}\sqrt{\rho_{\text{DM}}/2}$. Along the photon trajectory we take the eikonal propagation approximation $t\simeq z$ (with $c=1$).
The propagation equation represents a three-level driven system with period $T=2\pi/m_{a}$. It is convenient to remove the explicit time dependence induced by the axion background by transforming to a co-rotating frame that follows the axion-induced phase modulation. This is achieved by a time-dependent unitary transformation that eliminates the oscillatory off-diagonal term at the expense of introducing harmonic couplings between propagation eigenstates.
We perform a time-dependent unitary transformation $\Psi(z)=U(z)\Phi(z)$ to reallocate the rapid oscillations. Applying the rotation operator $U=u\oplus 1$, where $u$ acts on the photon subspace as
@@
the fast oscillating sinusoidal term is absorbed into a phase-dependent rotation angle $\theta(z)=\alpha\cos\phi(z)$ with $\alpha=\Delta_{\text{DM}}/m_{a}$ and $\phi(z)=m_{a}z+\varphi_{a}$. This rotation is chosen such that the term generated by $U^{\dagger}i\partial_{z}U$ exactly cancels the rapidly oscillating birefringence $\Delta_{F}$. Assuming negligible QED birefringence with $|\Delta_{\perp}-\Delta_{\parallel}|/2\ll m_{a}$ and $\Delta_{\mathrm{DM}}$, the effective Hamiltonian becomes:
@@
$\mathcal{H}^{\prime}(z)=-\begin{pmatrix}\Delta_{\gamma}&0&\Delta_{\mathrm{B}}\sin\theta\\ 0&\Delta_{\gamma}&\Delta_{\mathrm{B}}\cos\theta\\ \Delta_{\mathrm{B}}\sin\theta&\Delta_{\mathrm{B}}\cos\theta&\Delta_{a}\end{pmatrix}$, (8)
where $\Delta_{\gamma}=(\Delta_{\parallel}+\Delta_{\perp})/2$ is the mean photon refractive index. The unitary rotation is exact algebraically and does not change the quasienergy spectrum; it simply relocates the periodic dependence.
In the regime $\Delta_{\mathrm{B}}\ll|\Delta_{\gamma}-\Delta_{a}|$, the off-resonant photon mode can be adiabatically eliminated, allowing the dynamics to be analyzed by treating each photon mode-axion pair independently within the rotating-wave approximation. The resulting two-level systems are governed by the couplings $G^{(1)}=\Delta_{\mathrm{B}}\sin\theta$ and $G^{(2)}=\Delta_{\mathrm{B}}\cos\theta$. These couplings depend on $\theta(z)$; $\sin\theta$ and $\cos\theta$ can be written as linear combinations of $e^{\pm i\theta}$.
Using the Jacobi-Anger expansion, $e^{i\alpha\cos\phi}=\sum_{\ell}i^{\ell}J_{\ell}(\alpha)e^{i\ell\phi}$ with the Bessel function $J_{\ell}(\alpha)$, we expand $G^{(1,2)}$ into Fourier harmonics. The couplings can be written as
$G(z)=\sum_{\ell\in\mathbb{Z}}g_{\ell}e^{i\ell m_{a}z},$ (9)
where $g_{\ell}^{(1)}$ is nonzero only for odd $\ell$, while $g_{\ell}^{(2)}$ is nonzero only for even $\ell$. Explicitly, the $\ell$-th harmonic coupling scales as $g_{\ell}\sim\Delta_{\mathrm{B}}J_{\ell}(\alpha)$ up to phase factors, where the Bessel function $J_{\ell}(\alpha)$ behaves roughly as an oscillatory sine or cosine that decays proportionally to $\alpha^{-1/2}$.
In this co-rotating basis, the driven photon–axion system acquires a Floquet structure: the axion oscillation generates sidebands separated by integer multiples of $m_{a}$, and resonant mode mixing occurs when the propagation eigenvalue splitting matches one of these harmonics. In this basis, the axion-induced term acts as a periodic off-diagonal drive that couples propagation eigenstates separated in quasi-energy. Transitions between eigenstates become resonant whenever the energy mismatch $\Delta=\Delta_{\gamma}-\Delta_{a}$ is compensated by the absorption or emission of integer multiples of the driving frequency.
@@
We remove the free propagation phases by defining $\Phi=\operatorname{diag}\left(e^{i\Delta_{\gamma}z},e^{i\Delta_{a}z}\right)\tilde{\Phi}_{\gamma}$. Under this transformation, the $\ell$-th Fourier component of the coupling acquires a phase factor $e^{i(\Delta_{\gamma}-\Delta_{a}-\ell m_{a})z}$, with the corresponding detuning given by $\Delta_{\ell}=\Delta_{\gamma}-\Delta_{a}-\ell m_{a}$. Resonance occurs when the phase is stationary, yielding the Floquet resonance condition:
$\Delta_{\gamma}-\Delta_{a}\approx nm_{a},\quad n\in\mathbb{Z}.$ (10)
This condition characterizes coherent mode mixing within a time-periodic background, demonstrating that substantial conversion can occur even far from the standard level-crossing resonance $\Delta_{\gamma}=\Delta_{a}$.
Near a given resonance $n$, keeping only the corresponding Fourier mode amounts to a rotating-wave approximation. The two-state system undergoes Rabi oscillations with an effective frequency $\Omega_{\rm R}=\sqrt{|g_{n}|^{2}+(\Delta_{n}/2)^{2}}$. The conversion probability is given by
@@
The resonance is maintained as long as the detuning satisfies $|\Delta_{n}(z)|\lesssim 2|g_{n}|$, a regime where the peak conversion efficiency remains at least 50%. Note that the photon state $\gamma$ here refers to the effective photon mode under rotation. It can be verified that in the static limit $m_{a}\to 0$, all sidebands collapse and the system reduces exactly to conventional photon–axion mixing. Additionally, for small modulation amplitude $\Delta_{\mathrm{DM}}\ll m_{a}$, only the lowest harmonics contribute, and the effect reduces to a single-frequency perturbation of the standard oscillation probability.
More detailed derivations are provided in the supplementary material. An equivalent description in the circular polarization basis is also discussed in the supplementary material, where the axion background appears as a periodic modulation of the refractive indices. Related effects of axion-induced frequency modulation have been discussed previously, primarily in the context of birefringence or modifications of the dispersion relation [McDonald:2019wou]. In those treatments, the time-dependent phase is typically treated perturbatively or averaged over, and the associated Floquet sidebands are not explicitly identified. As a result, the resonant mode mixing and the corresponding Rabi oscillations discussed here were not recognized.
We emphasize that the Floquet resonance described here is fundamentally different from parametric instability or stimulated axion decay with $m_{a}\simeq 2\omega$. Here, the DM number density remains conserved, with the oscillating background providing the necessary energy ”kick” to bridge the momentum mismatch between the photon and axion dispersion relations. The Floquet coherent mixing in a periodic background is also dynamically distinct from the phenomenon of parametric instability. While both involve frequency matching, the former constitutes a unitary driven two-level problem characterized by stable mode-switching. In contrast, parametric instability requires a non-Hermitian structure for energy pumping (typically associated with sum-frequency channels  [^1] In the context of stimulated axion decay, the instability of the photon–axion system—considering both axion DM and magnetic fields but omitting the plasma background—is explored in Ref. [Masaki:2019ggg], where an instability condition $\sqrt{k_{\gamma}^{2}+m_{a}^{2}}+k_{\gamma}=nm_{a}$ is empirically identified.), which is kinematically inaccessible in the difference-frequency channel. Consequently, our resonance analysis leads to coherent oscillations that preserve the background number density, rather than the exponential growth characteristic of unstable systems. More discussion of this distinction is provided in the supplementary material.
@@
Representative Propagation.– To illustrate the physical implications of the Floquet resonance, we consider a representative photon propagation scenario in which the photon dispersion $\Delta_{\gamma}$ varies slowly along the propagation direction, while the axion background remains coherent over the relevant length scale. This setup captures the essential physics of resonance crossing without requiring detailed modeling of a specific astrophysical environment.
@@
As the photons propagate, the slowly varying dispersion causes the mismatch $\Delta(z)=\Delta_{\gamma}(z)-\Delta_{a}$ to sweep across the Floquet resonance condition $\Delta=nm_{a}$. When this occurs, the driven mode coupling becomes resonant, leading to efficient photon–axion conversion. Importantly, this resonance depends on temporal phase matching rather than spatial level crossing, and therefore persists in the absence of sharp gradients.
@@
Figure 1: Evolution of the photon–axion conversion probability $P_{\gamma\to a}$ along the jet of blazar. The blue solid line represents the case with an axion DM background, while the orange line shows the standard adiabatic conversion without DM. The vertical red dashed lines indicate the locations where the resonance condition $\Delta_{\gamma}(r)-\Delta_{a}\approx nm_{a}$ is satisfied.
Fig. 1 shows the photon–axion conversion probability $P_{\gamma\to a}$ as a function of propagation distance for a representative choice of parameters satisfying the Floquet resonance condition. A pronounced enhancement occurs when $\Delta(z)$ crosses $\Delta=m_{a}$, demonstrating coherent mode mixing induced by the axion-driven modulation. Away from resonance, the conversion probability remains small and oscillatory, consistent with conventional photon–axion mixing.
@@
The resonant region acts as a phase-matched mixing layer, in which the axion oscillation continuously supplies the energy mismatch required for conversion. The resulting behavior is analogous to Rabi oscillations in driven two-level systems, with the conversion probability bounded by unitarity.
@@
We have verified that the resonance persists under moderate variations of the dispersion profile and in the presence of stochastic fluctuations, provided that the axion field remains coherent over the resonance region. Thus, the effect is not a fine-tuned artifact of a specific background profile. Detailed numerical analysis is presented in the Supplementary Material.

### v2

@@ Abstract
Wave propagation in periodically time-dependent media can exhibit driven mode conversion that is absent in static or adiabatic descriptions. We show that photon propagation through a coherent axion dark matter background provides a natural realization of such driven dynamics. In the presence of a magnetic field, the oscillating axion field acts as a coherent temporal drive, inducing resonant photon–axion conversion when the mismatch between their dispersion relations is compensated by integer harmonics of the axion oscillation frequency, $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$ with $n\in\mathbb{Z}$. This driven resonance enables efficient mixing far from the conventional level-crossing regime and disappears entirely upon time averaging, explaining why it is missed in standard treatments. The process constitutes a unitary mode-conversion phenomenon that preserves the axion dark matter number density and is distinct from parametric instabilities or axion decay. A systematic description is naturally provided by Floquet theory. We develop a general framework for photon propagation in oscillating axion backgrounds and show that the resulting resonant mixing leads to characteristic polarization signatures, with potential implications for astrophysical observations such as blazar polarization.
@@
Introduction.– The propagation of waves in time-dependent media is a central problem across physics. When the properties of a medium vary periodically in time, wave propagation can exhibit driven mode mixing and resonant conversion phenomena that are absent in static or adiabatic descriptions. Such effects are well known in condensed matter and quantum optics [45, 41, 14, 32, 18], yet they have received little attention in the context of particle astrophysics.
@@
A prominent and well-motivated realization of a coherently oscillating medium arises if dark matter (DM) [50, 8] is composed of axions [34, 33, 48, 49, 13] or axion-like particles [36, 2, 16, 6, 21, 35, 27, 15, 3, 31]. In this case, the DM forms a classical field that oscillates at a frequency set by the axion mass over macroscopic coherence scales. Photons propagating through this background therefore experience a medium whose optical properties are intrinsically periodic in time.
Nevertheless, conventional treatments of photon–axion mixing [40, 39, 25] typically average over the axion oscillations or assume an effectively static background, thereby reducing the problem to a time-independent level-crossing or adiabatic conversion scenario. Most astrophysical studies consequently focus on birefringence [11, 12, 22] or Primakoff conversion [37, 46, 40] within quasi-static or adiabatic approximations.
@@
In this work, we show that these assumptions miss a qualitatively new class of resonant propagation effects. Photon–axion mixing in a coherent axion DM background is intrinsically a driven process, in which the temporal periodicity of the axion field acts as a coherent driver. Resonant mode conversion occurs when the mismatch between the photon and axion dispersion relations is compensated by integer harmonics of the axion oscillation frequency, $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$ ($n\in\mathbb{Z}$), allowing efficient conversion even far from the conventional level-crossing condition $\Delta_{\gamma}\approx\Delta_{a}$ [38, 24]. This resonance arises purely from temporal phase matching induced by the oscillating background and does not rely on spatial inhomogeneities or finely tuned environmental profiles. The effect relies on the coherence of the axion oscillations and disappears under time averaging, explaining why it is systematically missed in standard treatments. A systematic description of such periodically driven mixing is naturally provided by Floquet theory, originally formulated in Refs. [45, 41] and reviewed in Ref. [18].
@@
The driven resonance constitutes a unitary mode-conversion process, distinct from previously studied decay or instability phenomena [47, 23, 9, 10, 7], and leads to characteristic polarization signatures during propagation. While specific observational manifestations depend on environmental conditions, the underlying resonance is intrinsic to coherent axion DM and represents a previously unexplored regime of photon–axion mixing. More broadly, our results provide a general framework for wave propagation in time-dependent backgrounds and identify new resonant regimes beyond static or adiabatic descriptions.
@@
Driven photon–axion Mixing.– The axion may couple to electromagnetism through the operator $\mathcal{L}\supset-g_{a\gamma}aF\tilde{F}/4$, where $g_{a\gamma}$ is the photon–axion coupling with units of $\mathrm{GeV}^{-1}$, $a$ is the axion field, and $F$ is the electromagnetic field tensor. We consider photon propagation in the presence of a magnetic field and a coherently oscillating axion background (see also Refs. [19, 28, 29] for a similar system). The axion field is treated as a classical, homogeneous background,
@@
where $m_{a}$ is the axion mass, $\rho_{\text{DM}}$ is the DM density, and $\varphi_{a}$ an a priori unknown phase. The axion-photon mixing equations in the presence of an external magnetic field can be reduced to a system of first-order differential equations via the WKB approximation, assuming the beam energy $\omega$ is much larger than the axion mass $m_{a}$. We consider photon propagation along the $z$-axis in the presence of a transverse magnetic field $\mathbf{B}_{\mathrm{T}}$. In the linear polarization basis, the photon–axion system obeys the Schrödinger-like propagation equation,
@@
where the state vector $\Psi$ denotes the two photon polarization states and the axion field. The photon components correspond to linear polarizations oriented perpendicular and parallel to the direction of the external magnetic field. The dynamics is governed by a Hermitian mixing Hamiltonian with an explicit periodic time dependence induced by the oscillating axion background.
The static part $\mathcal{H}_{0}$ describes conventional photon–axion mixing in a magnetized medium and is given by [40, 39, 25]
@@
Throughout this work, all dispersion and mixing effects are parametrized in terms of $\Delta$-parameters following the standard photon–axion mixing notation. Here, $\Delta_{\perp,\parallel}$ represent the photon dispersion for the two linear polarizations, accounting for both plasma effects and vacuum birefringence. In the systems considered here, vacuum QED effects are included in $\Delta_{\perp,\parallel}$ but remain subdominant to the plasma contribution, $\Delta_{\perp,\parallel}\simeq-\omega_{\mathrm{pl}}^{2}/(2\omega)$. The effective plasma frequency $\omega_{\mathrm{pl}}=\sqrt{4\pi\alpha n_{e}/m_{e}}$ with $\alpha$ the fine-structure constant, where $n_{e}$ and $m_{e}$ denote the electron number density and mass, respectively. The axion mass-induced dispersion is given by $\Delta_{a}=-m_{a}^{2}/(2\omega)$, while the photon–axion mixing is governed by $\Delta_{\rm B}=g_{a\gamma}B_{\mathrm{T}}/2$.
The axion background generates an additional, explicitly time-dependent contribution
@@
arising from the axion-gradient-induced birefringence, analogous to but distinct from plasma Faraday rotation, with
@@
This axion-induced Faraday term is purely imaginary, reflecting its interpretation as a rotation generator. While similar time-dependent terms have appeared in Refs. [19, 28, 29], we formalize it here as a Floquet driving parameter $\Delta_{F}$ (see Supplemental Material for the full derivation from the Lagrangian level). This reformulation is not merely notational; it maps the propagation problem onto a quasi-energy framework.
Crucially, the driven contribution does not commute with the static mixing Hamiltonian, $\mathcal{H}_{0},\mathcal{H}_{1}(z)]\neq 0$. As a result, replacing the full time-ordered evolution operator by one generated by the averaged Hamiltonian is not a controlled approximation $\langle\mathcal{T}e^{-i\int\mathcal{H}(z)dz}\rangle\neq e^{-i\int\langle\mathcal{H}(z)\rangle dz}$. In particular, time averaging projects out the nonzero Floquet sectors responsible for the driven mode mixing, thereby eliminating the effect at the level of the Hamiltonian rather than suppressing it perturbatively.
The Hamiltonian is manifestly Hermitian, and its only explicit time dependence originates from the oscillating axion background through the birefringence term,
@@
where $\Delta_{\text{DM}}\equiv g_{a\gamma}\sqrt{\rho_{\text{DM}}/2}$ sets the amplitude of the axion-induced drive. This term oscillates at frequency $m_{a}$ and acts as a coherent temporal drive of the photon–axion system. Along the photon trajectory we adopt the eikonal approximation $t\simeq z$ (with $c=1$), appropriate for forward-propagating waves with a well-defined carrier frequency. In this limit, the temporal oscillation of the axion background translates into a periodic modulation along the propagation direction with period $T=2\pi/m_{a}$.
The resulting propagation equation describes a three-level system subject to coherent periodic driving. Rather than eliminating this time dependence, it is convenient to reallocate it by transforming to a co-rotating frame that follows the axion-induced phase modulation. This allows the driven nature of the mixing to be made explicit while preserving the quasienergy spectrum.
We therefore perform a time-dependent unitary transformation $\Psi(z)=U(z)\Phi(z)$ with $U=u\oplus 1$, where $u$ acts on the photon subspace as
@@
with $\sigma_{y}$ the second Pauli matrix. The rotation angle $\theta(z)=\alpha\cos\phi(z)$, with $\alpha=\Delta_{\rm DM}/m_{a}$ and $\phi(z)=m_{a}z+\varphi_{a}$, is chosen such that the contribution from $U^{\dagger}i\partial_{z}U$ exactly compensates the rapidly oscillating birefringence term $\Delta_{F}$. This transformation is algebraically exact and merely redistributes the periodic dependence among the Hamiltonian matrix elements.
In the following we neglect QED vacuum birefringence, assuming $|\Delta_{\perp}-\Delta_{\parallel}|/2\ll m_{a},\Delta_{\rm DM}$. This condition ensures that the two photon polarization states remain effectively degenerate on the scale of the axion-induced modulation, so that the driven mixing dynamics is not suppressed by polarization dephasing. We define $\Delta_{\gamma}\equiv(\Delta_{\perp}+\Delta_{\parallel})/2$ as the average photon dispersion. Then the effective Hamiltonian becomes
@@
$\mathcal{H}^{\prime}(z)=-\begin{pmatrix}\Delta_{\gamma}&0&\Delta_{\mathrm{B}}\sin\theta\\ 0&\Delta_{\gamma}&\Delta_{\mathrm{B}}\cos\theta\\ \Delta_{\mathrm{B}}\sin\theta&\Delta_{\mathrm{B}}\cos\theta&\Delta_{a}\end{pmatrix}$. (8)
The unitary transformation leaves the physical spectrum unchanged, but renders the periodic structure of the driven photon–axion mixing explicit, providing a natural starting point for the Floquet analysis below.
The full system is a three-level Floquet problem and can be treated as such. However, in the regime $\Delta_{\rm B}\ll|\Delta_{\gamma}-\Delta_{a}|$, one of the photon propagation eigenstates is far off resonance and can be adiabatically eliminated. The dynamics then reduces to two independent photon–axion subsystems, each of which can be analyzed within the rotating-wave approximation. The effective couplings are $G^{(1)}=\Delta_{\rm B}\sin\theta(z)$ and $G^{(2)}=\Delta_{\rm B}\cos\theta(z)$, acting as coherent time-dependent drives.
The axion-induced phase modulation is periodic, so the couplings can be expanded in a discrete set of harmonics with amplitudes controlled by Bessel functions $J_{\ell}(\alpha)$. Explicitly, this follows from the standard Jacobi–Anger expansion,
$e^{i\alpha\cos\phi}=\sum_{\ell\in\mathbb{Z}}i^{\ell}J_{\ell}(\alpha)e^{i\ell\phi}.$ (9)
Writing $\sin\theta$ and $\cos\theta$ in terms of $e^{\pm i\theta}$, the couplings can be decomposed into discrete Fourier harmonics,
$G(z)=\sum_{\ell\in\mathbb{Z}}g_{\ell}e^{i\ell m_{a}z}.$ (10)
Here $g_{\ell}^{(1)}$ ($g_{\ell}^{(2)}$) is nonzero only for odd (even) $\ell$, reflecting the parity structure of the two photon polarizations, and the magnitude of each harmonic scales as $g_{\ell}\sim\Delta_{\rm B}J_{\ell}(\alpha)$ up to phase factors.
In this co-rotating basis, the driven photon–axion system naturally acquires a Floquet structure [18]: the oscillating axion background generates a ladder of sidebands separated by integer multiples of $m_{a}$, and the periodic off-diagonal couplings mediate transitions between propagation eigenstates whose quasi-energies differ by $\ell m_{a}$.
@@
To make this explicit, we remove the free propagation phases and move to the interaction picture by defining $\Phi=\operatorname{diag}\left(e^{i\Delta_{\gamma}z},e^{i\Delta_{a}z}\right)\tilde{\Phi}$. The $\ell$-th Fourier component of the coupling then acquires a phase factor $e^{i\delta_{\ell}z}$, with detuning $\delta_{\ell}=\Delta_{\gamma}-\Delta_{a}-\ell m_{a}$. Resonant mode mixing occurs when this phase is stationary, yielding the Floquet resonance condition
$\Delta_{\gamma}-\Delta_{a}\approx nm_{a},\quad n\in\mathbb{Z},$ (11)
corresponding to vanishing detuning $\delta_{\ell}$ with the identification $\ell=n$. This condition generalizes the conventional level-crossing criterion to a periodically driven system, demonstrating that coherent photon–axion conversion can occur even far from $\Delta_{\gamma}=\Delta_{a}$. The axion background thus supplies discrete energy quanta that compensate the photon–axion dispersion mismatch, enabling coherent conversion in a regime inaccessible to static treatments.
Near a given Floquet resonance labeled by $n$, retaining only the corresponding Fourier component constitutes a rotating-wave approximation, valid when the nonresonant harmonics satisfy $|\delta_{\ell}|\gg|g_{\ell}|$ for $\ell\neq n$. The dynamics then reduces to a driven two-level system undergoing coherent Rabi oscillations [4, 44] with frequency
$\Omega_{\rm R}=\sqrt{|g_{n}|^{2}+(\delta_{n}/2)^{2}}.$ (12)
The photon–axion conversion probability takes the standard form
@@
where $\gamma$ denotes the effective photon propagation eigenstate in the co-rotating basis. Resonant conversion is sustained as long as the detuning remains within $|\delta_{n}(z)|\lesssim 2|g_{n}|$, for which the peak conversion probability exceeds $50\%$. This regime defines the finite bandwidth of each Floquet resonance.
@@
Several consistency limits are immediate. In the static limit $m_{a}\to 0$, all sidebands collapse and the system reduces exactly to conventional photon–axion mixing. For weak modulation, $\Delta_{\rm DM}\ll m_{a}$, only the lowest harmonics contribute, and the driven resonance reduces to a small, single-frequency perturbation of the standard oscillation probability.
Additional derivations are provided in the Supplemental Material, where we also discuss an equivalent formulation in the circular polarization basis. In that representation, the axion background appears as a periodic modulation of the refractive indices, providing an alternative but equivalent perspective on the driven mixing. Previous studies have considered axion-induced frequency modulation or birefringence [29], but typically treat the time dependence perturbatively or average over it, thereby missing the discrete Floquet sidebands and the associated resonant mode mixing.
Finally, we stress that the driven resonance identified here—sometimes loosely termed “parametric”—corresponds to a difference-frequency Floquet resonance and is dynamically distinct from parametric instability or stimulated axion decay at $m_{a}\simeq 2\omega$ [47, 23, 9, 10, 7]. In the present case the Hamiltonian remains Hermitian, the axion dark matter number density is conserved, and the oscillating background supplies only discrete energy quanta required to compensate the photon–axion dispersion mismatch. The result is bounded, unitary mode conversion characterized by coherent Rabi oscillations.
By contrast, parametric instability requires non-Hermitian energy pumping and access to sum-frequency channels, leading to exponential growth of field amplitudes. Such channels are kinematically inaccessible in the difference-frequency resonance considered here. The driven photon–axion mixing therefore realizes a stable Floquet resonance rather than an instability, as discussed further in the Supplemental Material.
@@
Representative Propagation.– To illustrate the physical consequences of Floquet resonances, we consider photon propagation in a medium with a slowly varying dispersion $\Delta_{\gamma}(z)$, in the presence of a coherent axion background over the relevant length scale. This setup isolates the driven resonance mechanism without reliance on detailed source modeling.
@@
As the photon propagates, the dispersion mismatch sweeps across the Floquet resonance condition $\Delta_{\gamma}(z)-\Delta_{a}\simeq nm_{a}$, activating resonant photon–axion mode conversion. Unlike conventional level-crossing scenarios, this resonance is governed by temporal phase matching and therefore does not require sharp spatial gradients or adiabaticity violation.
@@
Figure 1: Photon–axion conversion probability $P_{\gamma\to a}$ along a representative blazar jet. The blue line shows the case of driven Floquet mixing with oscillating DM background, while the orange line indicates standard adiabatic conversion with static DM background. Vertical red dashed lines mark locations where $\Delta_{\gamma}(r)-\Delta_{a}\approx nm_{a}$.
Figure 1 shows the resulting photon–axion conversion probability $P_{\gamma\to a}$ as a function of propagation distance for representative parameters. The driven Floquet case (blue) exhibits sharply enhanced conversion when the dispersion mismatch crosses a resonance band, whereas standard static mixing (orange) remains oscillatory and strongly suppressed. The vertical dashed lines indicate the locations where $\Delta_{\gamma}(z)-\Delta_{a}\simeq nm_{a}$.
@@
The resonant region acts as a phase-matched mixing layer in which the axion oscillation continuously compensates the photon–axion dispersion mismatch. The resulting dynamics is directly analogous to Rabi oscillations in a coherently driven two-level system, with conversion probabilities bounded by unitarity.
@@
We have verified that the resonance persists under moderate variations of $\Delta_{\gamma}(z)$ and in the presence of stochastic fluctuations, provided that the axion field remains coherent across the resonance region. The effect therefore does not rely on finely tuned environmental conditions. Further numerical details are presented in the Supplemental Material.